OpenStudy (anonymous):

Help me please. fan and medal Given the parent function f(x) = 2x, show a graph that shows f(x) + 1?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

all of my choices are curves

OpenStudy (anonymous):

Are you sure you've copied it exactly? It shouldn't be a curved graph.

hero (hero):

^What he said

OpenStudy (anonymous):

Given the parent function f(x) = 2x, which graph shows f(x) + 1?

OpenStudy (anonymous):

@Hero

hero (hero):

That's \(f(x) = 2^x\) which is completely different from \(f(x) = 2x\)

hero (hero):

\(f(x) = 2^x\) is an exponential function. \(f(x) = 2x\) is a linear function.

OpenStudy (anonymous):

ohhhhhhhhh.........

OpenStudy (anonymous):

what line to I look at?

hero (hero):

Hint: The graphs are color-coded

OpenStudy (anonymous):

A?

hero (hero):

Correct

OpenStudy (anonymous):

Using the graph below, calculate the average rate of change for f(x) from x = 0 to x = 2.

OpenStudy (anonymous):

x = −4 x = −2 x = 2 x = 4 would it be x=2?

OpenStudy (anonymous):

@Hero

OpenStudy (princeharryyy):

still need help @izthewiz7

OpenStudy (anonymous):

Yes, on the last thing two comments @princeharryyy

OpenStudy (princeharryyy):

do u kno wat average rate mean @izthewiz7

OpenStudy (anonymous):

the point at which two lines intersect? @princeharryyy

OpenStudy (princeharryyy):

nope @izthewiz7

OpenStudy (princeharryyy):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

I don't understand how I can get the answer from that

OpenStudy (princeharryyy):

can u find out what would be the f(x) from the given graph it would be

OpenStudy (princeharryyy):

@izthewiz7

OpenStudy (anonymous):

4? @princeharryyy

OpenStudy (princeharryyy):

see this curve can u guess wat type of function it would b @izthewiz7

OpenStudy (anonymous):

no. im just really confused.

OpenStudy (princeharryyy):

ohk tell me in wat u are confused ?

OpenStudy (anonymous):

everything.

OpenStudy (princeharryyy):

see this curve is f(x) = 3^x - 3

OpenStudy (princeharryyy):

so y2 occurs at x=2 => y2= 3^2-3 =>y2=6

OpenStudy (princeharryyy):

similarly y1 occurs at x=0 so, y1= 3^0 -3 =>-2

OpenStudy (princeharryyy):

so from the formula i gave u earlier (y2-y1)/(x2-x1) => (6-(-2))/(2-0) => 4 as your answer @izthewiz7

OpenStudy (princeharryyy):

r u getting anything @izthewiz7

OpenStudy (princeharryyy):

please b quick i don't have ample time @izthewiz7